OpenStudy (anonymous):

it is given that 2 is one of the roots of the quadratic equation x^2-5x+k=0. find the value of k.

OpenStudy (stamp):

(x - 2)(x + ?) = x^2 - 5x + k

OpenStudy (stamp):

-2*? = k

OpenStudy (anonymous):

\[x^2-5x+k=0\]now since 2 is one of the root so if we substitute x=2 in the equation the equation should be equal to 0 because 2 is a factor of it \[(2)^2-5*2+k=0\]\[4-10+k=0\]\[k=6\]

OpenStudy (anonymous):

thank you ^^

OpenStudy (anonymous):

my pleasure !